OpenStudy (anonymous):

Can someone explain Hooke's Law of Elasticity in the easiest and most basic terms ever?

OpenStudy (surry99):

tell me what your current understanding of Hooke's Law of Elasticity is?

OpenStudy (danjs):

When you pull on a spring, the restoring force is in the opposite direction. For an ideal spring, the magnitude of the force is proportional to the displacement. \[\left| F \right| \alpha \left| x \right|\] So if you assume a 1 dimensional problem, (no need for vectors), The force of the spring is: F=(-)some constant * x The properties of the materials defines the spring constant, k [N/m] F=-k*x Whenever this linear relation holds for the force and displacement, We call it Hook's Law.

OpenStudy (surry99):

I think you have most of it but here is who I think of it: Hooke's Law applies to linear springs. The "linear" part refers to two things: 1) The restoring force is linearly related to displacements ...F=-kx 2) The displacements must be small so that the spring returns to its original length. That is, the spring stiffness (k) does not change.

OpenStudy (danjs):

Right. If you look at a standard stress-strain curve, there is a linear part and then it curves concave down or up at the end. Hooke's Law holds in the first linear part of the graph and the spring constant is the slope of that line. When deformation happens, the spring will not return back to it's original length.

OpenStudy (surry99):

When permanent deformation occurs the spring will not go back to its original length because it has been stretched beyond the elastic range (linear portion of Stress-Strain Curve)